OpenStudy (anonymous):

billy has a box of coins that he uses when playing poker with friends the box currently contains 37 coins, consisting of pennies, dimes, and quarters. the number of pennies is equal to the number of dimes, and the total value is $2.23. how many of each denomination of coin does he have?

OpenStudy (saifoo.khan):

Salamz.

OpenStudy (anonymous):

heeyy salamz!

OpenStudy (anonymous):

the number of pennies which is equal to dimes is?.... and the number of quarters is?...

hero (hero):

Wow iraqichick. long time no see :D

OpenStudy (anonymous):

i know!! hiiii heroo!!

OpenStudy (anonymous):

been busy with new job and classes!! im finish up my final now!!

hero (hero):

I see

OpenStudy (pokemon23):

hero is awesome

OpenStudy (anonymous):

!

OpenStudy (anonymous):

yes he is!

OpenStudy (pokemon23):

I'm hero number 1 fan

hero (hero):

x + y + z = 37 .01x + .10y + .25z = 2.23

OpenStudy (anonymous):

how do i solve that

hero (hero):

by substitution

hero (hero):

First eliminate the z variable.

hero (hero):

z = 37 - x - y

hero (hero):

Now substitute that into the second equation where z is: .01x + .10y + .25(37-x-y) = 2.23

OpenStudy (anonymous):

oh god no lol

hero (hero):

Yeah, there's actually an easier way to do it. With this method, you end up substituting twice

hero (hero):

That's why nobody likes to do these